OpenStudy (anonymous):

Identify the series 2 + 5 + 8 + 11 + 14 + . . . as arithmetic, geometric or neither.

OpenStudy (anonymous):

@jim_thompson5910 @tcarroll010 @dpaInc @tcarroll010 @Hero @terenzreignz

terenzreignz (terenzreignz):

Okay, you know what a geometric sequence is?

OpenStudy (anonymous):

nope

terenzreignz (terenzreignz):

Right, for a sequence to be geometric, if you divide the second term by the first, you should get the same quotient as when you divide the third by the second, the fourth by the third, the fifth by the fourth, etc. So... what do you get when you divide the second term by the first?

OpenStudy (anonymous):

it is an arithhmetic series.becoz the common difference is 3 evry time....(d2-d1)=(d3-d2)=3